OpenStudy (anonymous):

find the exact value of sin(4pi/3-7pi/4)

OpenStudy (anonymous):

\[\sin (4\Pi -7\Pi/4)=\sin(9\Pi/4)=\sin(\Pi/4) becouse 9\Pi/4 and \Pi/4 are coterminal angles \]

OpenStudy (anonymous):

Therfore the answer is \[\sqrt{2/}2\]

OpenStudy (anonymous):

\[\sin\left(\frac{4\pi}{3}-\frac{7\pi}{4}\right)=\sin\frac{7\pi}{4}\cos\frac{4\pi}{3}-\sin\frac{4\pi}{3}\cos\frac{7\pi}{4}\]

OpenStudy (anonymous):

http://www.purplemath.com/modules/idents.htm